# Drinfeld rational fractions for affine Kac–Moody quantum symmetric pairs

### v2

@@ 9.5. Factorization theorem in higher rank
(9.19) $\eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \xi_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\boldsymbol{\phi}_{i}^{-}(-q^{-2}z^{-1})\boldsymbol{\phi}_{i}^{+}(-Cq^{-2}z){\mod U^{i}_{+}}[\negthinspace[z]\negthinspace],$
(9.20) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\quad\mod{(U_{-}^{i}\otimes U_{+}^{i}+U_{+}^{i}\otimes U_{-}^{i})}[\negthinspace[z]\negthinspace].$
@@
If $\mathbf{s}=0$, the corollary follows directly from Corollary 4.11 and Theorem 9.7. The case of non-standard parameters $\mathbf{s}\neq 0$ follows from a higher rank analogue of (2.24) and Theorem 7.5, using the fact that $\overline{\Delta}_{\mathbf{s}}(U_{\neq i,+})\subseteq{\mathbb{C}}\otimes U_{\neq i,+}$. The group-like property of the coproduct is deduced in the same way as in the proof of Theorem 7.5. ∎
Remark 9.9.
Although not necessary for our applications, it would be interesting to find out whether Corollary 9.8 can be strengthened so that $U_{+}^{i}$ and $U^{i}_{-}$ are replaced by $U_{+}$ and $U_{-}$, respectively. Preliminary calculations suggest that this is the case at least in rank two.

### v3

@@ 9.7. Factorization theorem in higher rank
(9.69) $\eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \xi_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz){\mod U_{+}}[\negthinspace[z]\negthinspace],$
(9.70) $\Delta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \eta_{\mathbf{s}}(\boldsymbol{\grave{\Theta}}_{i}(z))\otimes\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\quad\mod{(U_{-}\otimes U_{+}+U_{+}\otimes U_{-})}[\negthinspace[z]\negthinspace].$
@@
If $\mathbf{s}=0$, the corollary follows from Corollary 4.11 and Theorem 9.10. Let us explain this more precisely. In analogy to (3.4), the defining relations of $\mathcal{O}_{q}(\widehat{{\mathfrak{g}}})$ imply that
(9.71) $\boldsymbol{\grave{\Theta}}_{i}(z)=1+\frac{c_{i}^{-1}Cz^{2}\left(z^{-1}[A_{i,-1},\boldsymbol{A}_{i,+}(z)]_{q^{-2}}-q^{-2}[A_{i,0},\boldsymbol{A}_{i,+}(z)]_{q^{2}}\right)}{1-Cz^{2}}.$
By Proposition 9.10, we can write $\eta(\boldsymbol{A}_{i,+}(z))=\iota_{i}\eta(\boldsymbol{A}_{+}(z))+R(z)$ with $R(z)\in{U_{d_{i}\geq 1,+}}[\negthinspace[z]\negthinspace]$. Moreover, $\eta(A_{i,-1})=\iota_{i}\eta(A_{-1})+R_{-1}$, with $R_{-1}\in U_{d_{i}\geq 1,+}$, and $\eta(A_{i,0})=\iota_{i}\eta(A_{0})$. We call $R(z)$ and $R_{-1}$ “remainder terms”. Since the elements $\iota_{i}\eta(A_{-1})$ and $\iota_{i}\eta(A_{0})$ contain precisely homogeneous terms of degree ±$\pm\alpha_{i}$, multiplying them with remainder terms yields an element of $U_{+}$. Similarly, a product of remainder terms lies in $U_{+}$. It follows that
(9.72) $\eta(\boldsymbol{\grave{\Theta}}_{i}(z))\equiv$ $\ \iota_{i}\eta\left(1+\frac{c_{i}^{-1}Cz^{2}\left(z^{-1}[A_{-1},\boldsymbol{A}_{+}(z)]_{q^{-2}}-q^{-2}[A_{0},\boldsymbol{A}_{+}(z)]_{q^{2}}\right)}{1-Cz^{2}}\right)$
(9.73) $=$ $\ \iota_{i}\eta(\boldsymbol{\grave{\Theta}}(z))\equiv\iota_{i}(\boldsymbol{\phi}^{-}(z^{-1})\boldsymbol{\phi}^{+}(Cz))=\boldsymbol{\phi}_{i}^{-}(z^{-1})\boldsymbol{\phi}_{i}^{+}(Cz)$
modulo ${U_{+}}[\negthinspace[z]\negthinspace]$. In the third equivalence, Corollary 4.11 was used.
The case of non-standard parameters $\mathbf{s}\neq 0$ follows from a higher rank analogue of (2.24) and Theorem 7.5, using the fact that $\overline{\Delta}_{\mathbf{s}}(U_{+})\subseteq{\mathbb{C}}\otimes U_{+}$. The group-like property of the coproduct is deduced in the same way as in the proof of Theorem 7.5. ∎